# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling

### v2

@@ 3.2 Multi-Granularity Cross-modal Alignment
For capturing the local-level information, given that the street-view branch is expected to capture fine-grained details to enrich region embedding at ground level, approaches like [61, 30] rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens [76]. We leverage a fine-grained interaction mechanism to implement cross-modal alignment. Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective. Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens. The similar approach is also applied to text-to-image.
$\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(v_{ik_{1}}^{\top}t_{ik_{2}})}),\vspace{-0.5em}$ (4)
@@
where $v_{i}$ and $t_{j}$ denote the normalized embedding of street-view representation $\mathbf{z}_{I_{sv}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{sv}}$ in the $j$-th pair, respectively. The fine-grained token-level representation can be optimized through:
$\mathcal{L}_{CL}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}}).$ (6)
3.3 Pretraining & Fine-Tuning
Pretraining Stage. The overall objective of UrbanVLP can be defined as the simultaneous optimization of the aforementioned two losses:
$\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$ (7)
where $\alpha$ and $\beta$ are weight hyperparameter. Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
Fine-Tuning Stage. During the fine-tuning stage, we employ a linear probing approach [23]. This process begins by extracting frozen features $\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p}$ from the pretrained encoder for satellite images, street-view images, and street-view positions (The latter two are optional). Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{p})$. As depicted in Figure 2, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities. Note that in downstream tasks, text information is unnecessary as additional text generation time impedes real-time application.
4 Experiments
4.1 Experimental Setup
Dataset & Task Description.
In this paper, given the current lack of open-sourced datasets in the research community, we introduce a new benchmark dataset named $\mathtt{CityScape}$, which will be released upon paper notification. $\mathtt{CityScape}$ is unique in that it comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding high-quality textual descriptions. For the satellite image-text pairs, we adhere to [74] to utilize core area data of four cities in China. A comprehensive overview of $\mathtt{CityScape}$, along with relevant statistics, is provided in Appendix B.
Table 1: Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
Methods UrbanVLP StructuralUrban [41] Vision-LSTM [27] UrbanCLIP-SV [74] UrbanCLIP [74] PG-SimCLR [73] ViT [15]
@@ 4.1 Experimental Setup
Baselines for Comparison. We compare our method with seven recent baselines in the field of urban imagery-based socioeconomic prediction. The single granularity imagery-based urban region representation learning methods include: ViT [15], PG-SimCLR [73], UrbanCLIP [74], UrbanCLIP-SV [74]. We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM [27] and StructuralUrban  [41]. The in-depth overview of each baseline method is presented in the Appendix C.
@@
Evaluation Metrics. The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE). An increase in $R^{2}$, along with a decrease in RMSE and MAE values, signifies improved model accuracy. More implementation details are provided in Appendix D.
4.2 Performance Evaluation
Figure 4: $R^{2}$ results in Beijing.
@@ 4.2 Performance Evaluation
To further validate the effectiveness of the proposed UrbanVLP framework, we conduct comparisons with existing state-of-the-art methods on our proposed datasets. The comprehensive results are presented in Table 1, and we can draw the following conclusions: UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement. It can be seen that UrhanVLP outperforms the best baseline, StructuralUrban [41], by 3.3%, 2.3%, 2.5%, 4.9%, 1.0% and 2.4% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing. The similar performance improvement can also be seen in the dataset of the other three cities. Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing dataset are 1.8% and 2.1%, respectively. In order to provide a more intuitive display, we have created Radar Images in Figure 4 to visualize the performance of Beijing. The results further prove the effectiveness of our framework for urban region representation learning.
@@
Multi-granularity cross-modal alignment effectively incorporates fine-grained information, demonstrating superior performance. We can observe that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level. This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery. Concurrently, it is noticed that UrbanCLIP-SV underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation. This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
@@
Compared to other modalities, the textual modality facilitates a more explicit and comprehensive understanding of the region. In baseline models, there are works such as Vision-LSTM [27], which introduces trajectories, and PGSimCLR [73], which incorporates POI distributions to integrate external information and enhance regional representations. However, the superior performance of our UrbanVLP illustrates the efficacy of textual mobility as a high-information-density modality [23], which fully leverages the inherent knowledge of LLMs and explicitly enhances interpretability.
4.3 Ablation Studies
Figure 5: Ablation study on $\mathtt{CityScape}$-$\mathtt{Beijing}$.
@@ 4.3 Ablation Studies
We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch). More ablation experiments can be found in Appendix H.
The effectiveness of incorporating street-view branch. One of the main contributions of this work is the incorporation of multi-granularity information that reflects the urban spatial hierarchy. The street-view branch specifically represents micro-level modeling. It can be observed that the incorporation of street-view branch results in an average improvement of 3.56% in terms of $R^{2}$. The presence of the street-view branch introduces the capture of rich detailed information, thereby enhancing the precision of the modeling. Furthermore, it is also can be seen that while street-view imagery provides fine-grained clues, its individual impact is still overshadowed by that of satellite imagery. This aligns with the performance differential between UrbanCLIP and UrbanCLIP-SV discussed in the Section 4.2, underscoring the indispensable role of macro-level modeling
@@
The effectiveness of textual modality. In Table 1, we demonstrate the function of the textual modality by comparing UrbanVLP with a standard ViT-based model, which shares the same configuration as the unimodal visual encoder of UrbanVLP (encompassing both satellite and street-view branches). We then utilize the visual features extracted without textual augmentation to predict downstream urban indicators. As we can see, the lack of textual information leads to substantial performance degradation, underscoring the critical role of textual modalities in attaining a comprehensive visual representation. Similar findings were also reported in [74].
Exploration of the impact of location coordinates. The role of location coordinates has been underexplored in previous work of urban region representation learning. However, its importance has already been emphasized in work related to Geo-Localization [69, 79, 3]. In Figure 5, we evaluate the performance of the model without the Location Encoding branch. Although not as impactful as visual modalities, the inclusion of geospatial locations has yielded an average improvement of 1.5% in $R^{2}$. However, when relying solely on location information, the performance is suboptimal. This indicates that at the current stage, location data can only serve as an auxiliary enhancement and lacks the capacity for standalone modeling due to its deficiency in contextual semantic information.
Figure 6: Illustration of the quality of our generated description example.
Figure 7: Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
4.4 Qualitative Analysis
@@
We further investigate the quality of generated texts and the predictive performance of UrbanVLP in practice. Due to page limitations, more experimental analysis can be found in the Appendix.
@@
Illustration of the quality of our generated descriptions. To visually illustrate the quality of the generated description, we present an example in Figure 7. The generated description of the street-view image (a) is depicted in (b). Subsequently, we employ GPT4V [75] to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text. Furthermore, GPT4V also provides specific areas for improvement in (c). In (d), we utilize GPT4V to generate an image based on the description in (b). It is evident that the generated image bears a resemblance to the original one.
Figure 8: Our web platform. Zoom in or see Figure 17 for a better view.
Case Study for Predicted Results. Here we also show some predicted results in Figure 7 on $\mathtt{CityScape}$-$\mathtt{Beijing}$ dataset. As we can see, satellite images (a) and (b), although sharing similar layouts, differ significantly in their land use. (a) encompasses residential and campus areas, whereas (b) is indicative of industrial zones. Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators. The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions. Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated. The same principle applies to school and residential area as well. As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7(b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a). In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price. To showcase practical applications, we further develop a web-based system (Figure 8 and 17), with more details and discussion in Appendix I.
@@
Visualization of Region Representations. In this section, we map the region representations learned in UrbanVLP into a two-dimensional space using the PCA algorithm in Figure 10. This process aims to visualize how the semantic meanings of the regions are manifested within the high-dimensional representation space. As we can see, the three images on the left and on the right belong to different clusters, each exhibiting intra-cluster similarities and inter-cluster differences. Specifically, UrbanVLP effectively models regions into high-dimensional space, wherein satellite images with similar architectural layouts demonstrate spatial similarity. Consequently, this leads to a reasonable, accurate, and semantically rich representation of the regions.
@@
Figure 10: PCA visualization of urban imagery.
Figure 10 provides a further visualization with a population indicator, compared to UrbanCLIP. We use the dot color to indicate the population (log) in the corresponding regions of Beijing. As shown by the red dashed box, the regions modeled by UrbanVLP exhibit more significant clustering characteristics and greater distinction in high-dimensional space. Additionally, the lower parts of the two main band-shaped branches in the figure are also more distinctly separated. Points that are spatially close also have similar population values.
5 Conclusion and Future Work
@@
Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities. UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation. A text generation and calibration mechanism is also proposed to ensure high-quality description generation. It has achieved state-of-the-art results on the constructed $\mathtt{CityScape}$ and downstream dataset. In the future, the further research direction can be two-fold: 1) Adding additional modalities such as POIs to enhance information richness, and 2) enabling the design of better model architectures such as MAE [23] to efficiently utilize existing data source.
Appendix / supplemental material
Appendix A Formulation

### v3

@@ Introduction
The street-view branch aims to enrich region embedding with fine-grained local-level information. However, previous approaches like (Radford et al. 2021; Jia et al. 2021) rely exclusively on global feature similarity within each modality, disregarding the necessity for fine-grained alignment, such as the correspondence between visual objects and textual tokens (Yao et al. 2022). To address this challenge, we leverage a fine-grained interaction mechanism to implement cross-modal alignment at the local level. Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective. We first compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens. The similar approach is also applied to text-to-image process.
$\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(v_{ik_{1}}^{\top}t_{ik_{2}})}),$ (6)
@@
where $v_{i}$ and $t_{j}$ denote the normalized embedding of street-view representation $\mathbf{z}_{I_{sv}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{sv}}$ in the $j$-th pair, respectively. The fine-grained token-level representation can be optimized via:
$\mathcal{L}_{CL}=-\frac{1}{N}\Bigg{(}$ $\sum_{i}^{N}\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}$ (8)
$+\sum_{i}^{N}\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}\Bigg{)}.$
Methods UrbanVLP StructuralUrban Vision-LSTM UrbanCLIP-SV UrbanCLIP PG-SimCLR ViT
@@
Table 1: Socioeconomic indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
Pretraining & Fine-Tuning
Pretraining Stage. The overall objective of UrbanVLP can be defined as the joint optimization of the above two losses:
$\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$ (9)
where $\alpha$ and $\beta$ are hyperparameters for a trade-off. Through backpropagation optimization, we achieve multi-granularity cross-modal alignment, resulting in robust encoders.
Fine-Tuning Stage. As depicted in Figure 3 Right, during the fine-tuning stage, we employ a linear probing approach (He et al. 2022) which begins by extracting $\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p}$ features from the pretrained encoder for satellite images, street-view images, and street-view positions. Subsequently, these features are fused together, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p})$. It is noteworthy that in downstream tasks, textual information is unnecessary since the knowledge embedded in the text has already been imparted to other modality encoders through pretraining, and additional text generation time hinders real-time application.
Experiments
Experimental Setup
Dataset & Task Description. Given the current lack of open-sourced datasets in the research community, we introduce a new benchmark dataset named $\mathtt{CityView}$, which will be released upon paper notification. $\mathtt{CityView}$ is unique in that it comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding high-quality textual descriptions. We adhere to (Yan et al. 2024) to cover core area data of four cities in China. A comprehensive overview of $\mathtt{CityView}$, along with relevant statistics, is provided in Appendix C.
Baselines. Following the established practice (Yan et al. 2024; Xi et al. 2022; Liu et al. 2023d) in this area, we compare our method with seven recent baselines in the field of imagery-based USI prediction. The single-granularity urban imagery-based methods include:
ViT (Dosovitskiy et al. 2021), PG-SimCLR (Xi et al. 2022), UrbanCLIP (Yan et al. 2024), UrbanCLIP-SV (Yan et al. 2024).
We also apply multiple baselines for multi-granularity urban imagery-based USI prediction: Vision-LSTM (Huang et al. 2023) and StructuralUrban (Li et al. 2022c). They are all trained using data from our $\mathtt{CityView}$ dataset. The in-depth overview of baselines is presented in Appendix E.
@@
Evaluation Metrics. We follow (Yan et al. 2024) to evaluate our method in terms of: the coefficient of determination $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE). An increase in $R^{2}$, along with a decrease in RMSE and MAE values, signifies improved model accuracy.
Implementation Details. The implementation details are provided in Appendix F.
Performance Evaluation
To evaluate our UrbanVLP framework, we conduct comparisons with existing state-of-the-art methods on our proposed $\mathtt{CityView}$ datasets. Table 1 presents the overall results, from which we can obtain the following findings:
@@
1) UrbanVLP significantly outperforms the baselines which employ single/multi-granularity imagery. It can be seen that UrbanVLP surpasses the best baseline (StructuralUrban) by 3.3%, 2.3%, 2.5%, 4.9%, 1.0% and 2.4% in terms of $R^{2}$ for all six indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing. Similarly, we can witness consistent improvements in the other three cities. In addition, the average reduction of UrbanVLP in terms of RMSE and MAE on the Beijing dataset are 1.8% and 2.1%, respectively. For a more intuitive display, we visualize the overall performance on Beijing and Shenzhen in Figure 2, where the results further prove the versatility of our framework for USI prediction.
@@
2) Integrating multi-granularity imagery can lead to superior performance against baselines. This merit can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery. Meanwhile, UrbanCLIP-SV underperforms UrbanCLIP, revealing that street-view imagery (while rich in details) lacks a macro view, resulting in suboptimal outcomes when used in isolation. This also underscores the irreplaceability of satellite imagery serving as a macro visual modality.
@@
3) Compared to other modalities (e.g., POIs), the textual modality facilitates a more comprehensive understanding of the region. We include various baselines for comparison, such as PGSimCLR which incorporates POI distributions to integrate external information and enhance regional representations. Despite their promising results, our UrbanVLP illustrates the efficacy of integrating textual data as a information-compact modality (He et al. 2022), which fully leverages the inherent knowledge of LLMs and explicitly enhances interpretability in the training phase.
Ablation Studies
@@
As shown in Figure 5, we conduct ablation studies to examine each component in UrbanVLP on the $\mathtt{CityView}$-Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch). More discussion can be found in Appendix J.
Figure 5: Ablation study on $\mathtt{CityView}$-$\mathtt{Beijing}$.
Effects of Street-View Branch. One of our contributions is the incorporation of multi-granularity information that reflects the urban spatial hierarchy. It can be observed that the incorporation of street-view branch results in an average improvement of 3.56% in terms of $R^{2}$. This enhancement is attributed to the street-view branch’s ability to capture rich, detailed information, thereby improving modeling precision.
@@
Effects of Textual Modality. In Table 1, we demonstrate the function of the textual modality by comparing UrbanVLP with a standard ViT-based model, which shares the same configuration as the unimodal visual encoder of UrbanVLP. We then utilize the visual features extracted without textual augmentation to predict downstream USI. As we can see, the lack of textual information leads to substantial performance degradation, underscoring the critical role of textual modalities in attaining a comprehensive visual representation. Similar findings were also reported in (Yan et al. 2024).
Effects of Location Coordinates. In Figure 5, we evaluate the performance of the model without the Location Encoding branch. Although not as impactful as visual modalities, the inclusion of geospatial locations has yielded an average improvement of 1.5% in $R^{2}$. It is noteworthy that we utilize a well-pretrained location encoder and freeze it within our framework. Consequently, its generalizability is ensured, preventing overfitting to specific cities in $\mathtt{CityView}$ dataset.
Qualitative Analysis
@@
We further investigate the quality of generated texts and the predictive performance of UrbanVLP in practice. More empirical analysis can be found in Appendix J.
@@
Illustration of the quality of our generated descriptions. To visually illustrate the quality of the generated description, we present an example in Figure 6. The generated description of the street-view image (a) is depicted in (b). Subsequently, we employ GPT4V (Yang et al. 2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text. Furthermore, GPT4V also provides specific areas for improvement in (c). In (d), we utilize GPT4V to generate an image based on the description in (b). It is evident that the generated image bears a resemblance to the original one.
Figure 6: An example of generated text descriptions.
Case Study for Predicted Results. Here we also show some predicted results in Figure 7 on $\mathtt{CityView}$-$\mathtt{Beijing}$ dataset. Satellite images (a) and (b), though similar in layout, differ significantly in land use. (a) encompasses residential and campus areas, whereas (b) represents industrial zones, leading to entirely different socioeconomic characteristics. The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial parks, whereas the disparities in population are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions.
As observed, UrbanCLIP’s predictions fail to effectively differentiate between the two, as the downstream metrics in Figure 7(b) still lean towards predicting a homogenized socioeconomic area similar to (a). In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socioeconomic attributes of the two in terms of carbon emission and GDP.
Figure 7: Case study of UrbanCLIP (UC) and our UrbanVLP (UV).
Visualization of Region Representations
@@
In this section, we map the region representations learned in UrbanVLP into a two-dimensional space using the PCA algorithm in Figure 8. As we can see, the three images on the left and on the right belong to different clusters, each exhibiting intra-cluster similarities and inter-cluster differences. Therefore, UrbanVLP could effectively model regions into high-dimensional space, wherein satellite images with similar architectural layouts demonstrate spatial similarity. Our framework learns a reasonable, accurate, and semantically rich region representation for USI prediction.
@@
Figure 9: Transferability test on $R^{2}$ between UrbanVLP and UrbanCLIP, on the Carbon indicator across 4 cities.
Transferability Study
In this study, we explore the practical application of USI prediction within a transfer learning context (Park et al. 2022; Liu et al. 2023d). Specifically, We use visual encoders trained on source city data and fine-tune them on target city data. We experiment with different pairs of source and target cities and present the carbon emission prediction in Figure 9. As we can see, on 16 source-target city pairs, UrbanVLP achieves an average $R^{2}$ of 0.574, while that of UrbanCLIP is 0.495. Additionally, the $R^{2}$ values show greater similarity between Beijing and Shanghai, as well as between Shenzhen and Guangzhou. This observation aligns with the geographical proximity of Beijing to Shanghai and Guangzhou to Shenzhen. These findings confirm the robust transferability of our UrbanVLP model in urban areas, despite the previously mentioned differences in geological and demographic characteristics among the selected cities.
Analysis about Hyperparameters
In Figure 10 (a), we explore different combinations of hyperparameter values for loss combination. It can be seen that the settings we choose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results. We also study the maximum value of street-view images within one region during training, which is set as 25 in our paper. As is shown in Figure 10 (b), the $R^{2}$ value steadily increases with the number of street-view images, peaking at 25. Therefore, our setting is the optimal.
Figure 10: (a) Hyperparameter for loss combination. (b) Experiments for the maximum number of street-view images.
Practicality
To showcase the practical applications, we further develop a web-based system (Figure 19 in Appendix), with more details and discussion in Appendix K. This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. It enhances social planning and sustainable development by providing a data-driven platform that deepens understanding of socioeconomic indicators.
Conclusion and Future Work
@@
USI prediction plays a significant role for understanding societal patterns and dynamics. UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation. A text generation and calibration mechanism is also proposed to ensure high-quality description generation. It has achieved state-of-the-art results on the constructed $\mathtt{CityView}$ dataset. Future research could incorporate additional modalities such as POIs, road networks, and building footprints to enhance information richness and introduce a broader perspective.
Acknowledgments
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
A    Formulation